OpenStudy (anonymous):

Write the following equation in slope-intercept form. - y + 6 = 0

OpenStudy (anonymous):

Slope intercept form involves getting y by itself. So you have y = mx + b. So start by subtracting the 6.

OpenStudy (anonymous):

the question is wrong. it's x/2-y+6=0

OpenStudy (anonymous):

\[\frac{x}{2} - y + 6 = 0\] ?

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

Ok, well to get y by itself, all you have to do is literally add it to the other side.

OpenStudy (anonymous):

x/2+6=y+0?

OpenStudy (anonymous):

Yes. y + 0 is simply "y"

OpenStudy (anonymous):

clearly..then you subract 6 which is x/2=y-6 then you cant even get y alone???

OpenStudy (anonymous):

No, your solution is as it was. That is slope intercept form: \[y = \frac{1}{2}x + 6\]

OpenStudy (anonymous):

Remember that: \[\frac{1}{2} x = \frac{x}{2}\]

OpenStudy (anonymous):

would it just be y=x/2+6?

OpenStudy (anonymous):

Yep.